A market surveyor wishes to know how many energy drinks teenagers drink each week. They want to construct a 95% confidence interval for the mean and are assuming that the population standard deviation for the number of energy drinks consumed each week is 0.6. The study found that for a sample of 233 teenagers the mean number of energy drinks consumed per week is 4.1. Construct the desired confidence interval. Round your answers to one decimal place.
Answer #1

Solution:

Confidence Interval for population mean using z distribution

Given,

\bar x = 4.1 ....... Sample mean

\sigma   = 0.6 ........Population standard deviation

n = 233 ....... Sample size

Note that, Population standard deviation(\sigma) is known..So we use z distribution.

Our aim is to construct 95% confidence interval.

\therefore c = 0.95

\therefore\alpha = 1- c = 1- 0.95 = 0.05

\therefore  \alpha/2 = 0.05 \slash 2 = 0.025 and 1- \alpha /2 = 0.975

Search the probability 0.975 in the Z table and see corresponding z value

\thereforeZ_{\alpha/2} = 1.96 for the 95% Confidence interval

Now , confidence interval for mean(\mu) is given by:

\bar x - Margin. of .error < \mu < \bar x + Margin. of .error

  \bar x-Z_{\alpha/2}*(\sigma/\sqrt{n}) < \mu < \bar x+Z_{\alpha/2}*(\sigma/\sqrt{n})

4.1 -1.96*(0.6/ \sqrt{} 233)   < \mu < 4.1 +1.96*(0.6/ \sqrt{} 233)

4.1 - 0.07704232157 < \mu <   4.1 + 0.07704232157

4.0 < \mu < 4.2

is the required 95% confidence interval for mean....

Answer is (4.0,4.2)